OpenStudy (anonymous):

What is the solution to 89 + k = –42?

OpenStudy (anonymous):

subtract 89 from -42

OpenStudy (anonymous):

subtract -41 to 89 and you will find the value ok k..

OpenStudy (anonymous):

89 + k = -42 -89 -89 k = -131 easy :)

OpenStudy (anonymous):

k=-131